OpenStudy (anonymous):

Rewrite in simplest rational exponent form √x * 4√x

OpenStudy (anonymous):

@welshfella @Nnesha Do you understand this?

Nnesha (nnesha):

you can convert root to an exponent form \[\huge\rm \sqrt[n]{x^m} =x^\frac{ m }{ n }\] index will become the denominator of the fraction so you can convert square root to 1/2 exponent

OpenStudy (anonymous):

4^1/2?

Nnesha (nnesha):

no 4 isn't under the square root \[\huge\rm \sqrt{x} \times 4\sqrt{x}\] this is ur question so convert sqrt{x} to exponent form

OpenStudy (anonymous):

x^1/2?

Nnesha (nnesha):

yes right \[\huge\rm x^\frac{ 1 }{ 2 } \times 4x^\frac{ 1 }{ 2 }\] now apply the exponent rule when we multiply same bases we should `add` exponents \[\huge\rm x^m \times x^n=x^{m+n}\]

OpenStudy (anonymous):

4x^2/4?

Nnesha (nnesha):

hmm \[\huge\rm 4x^{\frac{ 1 }{ 2 }+\frac{ 1 }{ 2 }}\] 1/2+1/2= ???

OpenStudy (anonymous):

1?

Nnesha (nnesha):

yes right we should find the common denominator \[\frac{ 1 }{ 2 }+\frac{ 1 }{ 2 }=\frac{ 1+1 }{ 2 }\] denominator are the same so just ~ADD~ the number in the numerator

OpenStudy (anonymous):

2/2

OpenStudy (anonymous):

4x^2/2

Nnesha (nnesha):

yes right 2/2 = ?

OpenStudy (anonymous):

Do you divide that?

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

So, 1?

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

So would the answer just be 4x?

Nnesha (nnesha):

yes right!

OpenStudy (anonymous):

Thank you for your help

Nnesha (nnesha):

np good work! :=)